40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Jun 1 13:56:52 2009, Alberto, Update, PSL, Laser Power after fixing the laser chiller MOPAtrend.jpg
    Reply  Mon Jun 1 15:01:31 2009, rana, Update, PSL, Laser Power after fixing the laser chiller: more traces Picture_7.png
       Reply  Mon Jun 1 19:19:26 2009, rana, Update, PSL, 1000 days of hour-trend u.png
Message ID: 1639     Entry time: Mon Jun 1 15:01:31 2009     In reply to: 1636     Reply to this: 1640
Author: rana 
Type: Update 
Category: PSL 
Subject: Laser Power after fixing the laser chiller: more traces 
If you look at the correlation between RMTEMP and HTEMP, you see what we knew: namely that there
was a 1:1 correlation before. After the chiller fix, I can see no correlation between the room and
amplifier temperature at the resolution of 10:1. So the chiller loop has a gain > 10 at 24 hour time
scales.

I don't understand why the PMC looks more stable.
Attachment 1: Picture_7.png  85 kB  | Hide | Hide all
Picture_7.png
ELOG V3.1.3-